OpenStudy (anonymous):

Using the definition of the derivative, differentiate the function f(x) = x + 9/x

OpenStudy (kc_kennylau):

you mean using first principle?

OpenStudy (anonymous):

\[f'(x)=\lim_{h\to0}\frac{f(x+h)-f(x)}{h}\] Not sure what your given function is, could be \[f(x)=x+\frac{9}{x}~~\text{ or }~~f(x)=\frac{x+9}{x}\]

OpenStudy (kc_kennylau):

no matter which case, just try to plug it in :)

OpenStudy (anonymous):

it's the first function you wrote SithsAndGiggles

OpenStudy (anonymous):

\[\lim_{h\to0}\frac{\left(x+h+\frac{9}{x+h}\right)-\left(x+\frac{9}{x}\right)}{h}\\ \lim_{h\to0}\frac{h+\frac{9}{x+h}-\frac{9}{x}}{h}\\ \lim_{h\to0}\frac{h}{h}+\lim_{h\to0}\frac{\frac{9}{x+h}-\frac{9}{x}}{h}\\ \lim_{h\to0}\frac{h}{h}+\lim_{h\to0}\frac{\frac{9x-9(x-h)}{x(x+h)}}{h}\\ \lim_{h\to0}\frac{h}{h}+\lim_{h\to0}\frac{\frac{9h}{x(x+h)}}{h}\\ \lim_{h\to0}\frac{h}{h}+\lim_{h\to0}\frac{9}{x(x+h)}\\ ~~~~~~~~~~~~~~~~~~~~\vdots \]

OpenStudy (anonymous):

Missing a minus sign on the last term: \[~~~~~~~~~~~~~~~~~~~~\vdots\\ \lim_{h\to0}\frac{h}{h}\color{red}{-}\lim_{h\to0}\frac{9}{x(x+h)}\\ ~~~~~~~~~~~~~~~~~~~~\vdots \]